A $32,000 loan at 12% compounded quarterly is to be repaid by equal quarterly payments over a seven-year term.

a.

What will be the principal component of the sixth payment? (Round your answer to 2 decimal places.)

b.

What will be the interest portion of the twenty-second payment? (Round your answer to 2 decimal places.)

c.

How much will the loan’s balance be reduced by Payments 10 to 15 inclusive? (Round your answer to 2 decimal places.)

d.

How much interest will be paid in the second year? (Round your answer to 2 decimal places.)
Answer #1

SEE THE IMAGE. ANY DOUBTS, FEEL FREE TO ASK. THUMBS UP PLEASE

SOLVED WITH BA II PLUS CALCULATOR

SET CALCULATOR : C/Y= 4 AND P/Y = 4

C w ENG 06:54 01-08-2020 70 35 Х E1967 . A B с D E F I G J K set C/Y = P/Y = 4 BA II PLUS CALCULATOR 1967 1968 1969 1970 1971

C W ENG 06:54 01-08-2020 70 35 Х G2010 . A B с D E F. G Н J K LOAN BALANCE AFTER 21 PAYMENTS N = 7 [28-21] 1/Y = 12.00 PMT =

C W ENG 06:54 01-08-2020 70 35 Х G2016 . A B с D E F. G I J K 2016 2017 2018 2019 2020 2021 2022 2023 2024 2025 LOAN BALANCE